OpenStudy (awkwardpanda):

Solve inequality. Graph solution. 1. 57-4t greater than or equal to 13.

OpenStudy (nirmalnema):

57-4t >= 13 57-13>=4t t <=11

OpenStudy (awkwardpanda):

Thank you, but how did you get 11?

OpenStudy (anonymous):

57-4t>=13 -4t>=13-57 -4t>=-44 to remove both negatives, change sign 4t=<44 t=<11 Hope this can help!

OpenStudy (awkwardpanda):

Thank you! can you help with another please? \[-18-5y \ge52\]

OpenStudy (anonymous):

-18-5y>=52 -5>=70 y>=14

OpenStudy (anonymous):

i mean y>=-14 sorry

OpenStudy (awkwardpanda):

Thank you! SO on the graph i would shade everything on the left side or right?

OpenStudy (anonymous):

really sorry I'm still not up to that :)